A hydraulic press has a RAM of 30 cm diameter and plunger of 5 cm diameter . Find the weight lifted by the hydraulic press when the force applied at the plunger is 400N
1
Expert's answer
2021-07-22T16:06:02-0400

Since the pressure are same,

Thus (F/A)=f/a

F=fA/a

F=(400x30^2/5^2)

F=400x36=14400N